Current Attempt in Progress During January 2022, Blue Spruce Corp.paid a cash dividends of $2320. This transaction O reducesDuring February 2022, its first month of operations, the owner of Metlock, Inc. invested cash of $90000. Metlock had cash salAt September 1, 2022, Metlock, Inc. reported a cash balance of $126000. During the month, Metlock collected cash of $54000 anDuring 2022, its first year of operations, Splish Brothers Inc. had revenues of $156000 and expenses of $79200. The businessCurrent Attempt in Progress Sheridan Company compiled the following financial information as of December 31, 2022: Service re
Answer #1

1.

Cash dividend paid = $2,320

This transaction would decrease stockholders' equity and assets by $2,320

First option is correct.

2.

Owner investment = $90,000

Cash sales = $18,000

Expenses paid = $31,500

Ending cash balance = Owner investment + Cash sales - Expenses paid

= 90,000 + 18,000 - 31,500

= $76,500

Second option is correct.

3.

Beginning cash = $126,000

Cash collection = $54,000

Cash disbursements = $90,000

Ending cash balance = Beginning cash + Cash collection - Cash disbursements

= 126,000 + 54,000 - 90,000

= $90,000

First option is correct.

4.

Revenues = $156,000

Expenses = $79,200

Dividend = $43,200

Net income = Revenues - Expenses

= 156,000 - 79,200

= $76,800

Retained earnings = Net income - Dividend

= 76,800 - 43,200

= $33,600

Third option is correct.

5.

Total assets = Equipment + Cash + Supplies + Accounts receivables

= 234,000 + 204,000 + 28,000 + 90,000

= $556,000

First option is correct.